Solve this equation.
a squared + 15 squared = 17 squared

Answer by mukhopadhyay(490)

a squared + 15 squared = 17 squared
=>a^2+15^2=17^2
=>a^2=17^2-15^2
=>a^2=(17+15)(17-15)
=>a^2=32*2
=>a^2=64
=>a=8 or a=-8